# TRACE: Time SeRies PArameter EffiCient FinE-tuning

### v2

@@ Abstract
We propose an efficient fine-tuning method for time series foundation models, termed TRACE: Time Series Parameter Efficient Fine-tuning. While pretrained time series foundation models are gaining popularity, they face the following challenges: (1) Unlike natural language tasks, time series data vary in frequency, channel numbers, historical/prediction lengths. For long-term forecasting tasks in particular, tailored fine-tuning can significantly enhance performance.(2) Existing parameter-efficient tuning methods like LoRA remain applicable but require adaptation to temporal characteristics.
To address these challenges, our TRACE framework introduces two key innovations: (1) Gated DSIC (Gated Dynamic Simulation Importance Calculation), an unbiased LoRA module importance selection mechanism that ensures conditional parameter consistency before and after masking. Experiments demonstrate that Gated DSIC outperforms common fine-tuning. (2) Reconstructed prediction heads for long-term forecasting tasks, which achieve comparable or superior performance to linear probing heads while drastically reducing parameter counts.
@@ 1 Introduction
Background: Accurate time series forecasting and anomaly detection are crucial research areas. Time series modeling has evolved significantly, from traditional Exponential Smoothing[1] and ARIMA[2] models to deep learning-based Transformers. Recently, pre-trained time series backbone models have gained attention, improving forecasting accuracy and enabling new solutions for anomaly detection. Inspired by pre-training strategies in natural language processing (NLP), these models leverage cross-domain technologies, broadening their applications across fields.
Current Methods: Classical deep learning models like LSTMs [3] and GRUs [4] require task-specific training, which is computationally expensive and struggles with cross-dataset generalization. Emerging pre-trained time series models address this by capturing general temporal patterns through unsupervised learning and adapting to downstream tasks via fine-tuning. However, current methods largely rely on either full parameter fine-tuning or linear probing.
Full parameter fine-tuning adjusts all model parameters but incurs high computational costs, making it impractical for resource-constrained settings. It also underutilizes the model’s pre-trained generalization capabilities. Linear probing, which fine-tunes only the final layer, is computationally cheaper but fails to fully leverage the rich semantic information in intermediate layers. For long-term forecasting, linear probing introduces excessive parameters, increasing overfitting risks and inefficiency.
These limitations highlight the need for more efficient transfer learning strategies and improved fine-tuning methods for time series tasks, particularly in long-term forecasting.
Our solution: To address the issue of excessive parameter size in the prediction heads of long-term forecasting tasks, we propose a novel solution of restructuring the prediction head. Our approach reduces the parameter count by more than 70% while simultaneously improving performance. By reducing the model’s complexity, we not only alleviate the risk of overfitting but also effectively leverage the generalization capability of the base model, making it more adaptable to downstream tasks.
Additionally, we introduce an efficient fine-tuning method for time series pretraining models, called Gated DSIC(Gated Dynamic Simulation Importance Calculation) Low-Rank Adaptation. Gated DSIC employs LoRA Gates to control each LoRA module individually, allowing for precise calculation of module importance. By eliminating biases in the importance computation, our method accurately captures the contribution of each LoRA module. This fine-tuning technique enables efficient adjustment of large-scale pre-trained models, even on consumer-grade GPUs, without sacrificing performance.
Contributions: This paper introduces the TRACE(Time Series Parameter Efficient Fine-tuning) method, an unbiased selection technique for LoRA modules. It mainly consists of the following four parts:
[…]
  • Gated DSIC LoRA: This introduces a gating mechanism for LoRA modules, enabling parameter importance calculation and masking during fine-tuning. Gated DSIC ensures unbiased measurement of each LoRA module’s contribution, improving fine-tuning accuracy.
  • We conducted comparative experiments on a diverse set of temporal tasks and observed the following: our method achieves superior performance with fewer parameters in the prediction head reconstruction compared to linear probing. Additionally, when applying LoRA fine-tuning to the backbone Transformer, Gated-DISC outperforms full LoRA and AdaLoRA using fewer LoRA modules.
  • In addition to time series tasks, our method was evaluated against various LoRA approaches on the GLUE benchmark and three challenging question-answering datasets, demonstrating its effectiveness in natural language tasks and providing valuable insights for future research.
@@ 2.1 Time Series Modeling
Statistical Methods: Early methods like Moving Average (MA) smooth trends but fail to capture complex dynamics. ARIMA [2] improves accuracy by integrating Autoregressive (AR) and MA components but struggles with nonlinear relationships. VAR [5] models multivariate relationships but suffers from dimensionality issues. These methods are limited to univariate modeling and cannot effectively use covariates.
Machine Learning Methods: Gradient boosting trees (e.g., XGBoost [6], LightGBM [7]) leverage feature engineering strategies, such as lag features and sliding window statistics, to model nonlinear patterns. While outperforming statistical methods, their performance depends heavily on feature quality and fixed window lengths limit long-term dependency modeling.
Deep Learning Methods: RNNs and variants like LSTM [3] and GRU [4] capture temporal dependencies but suffer from inefficiency and vanishing gradients for long-range dependencies. TCNs [8] expand receptive fields via dilated convolutions but are constrained by fixed kernels.
Transformer Architectures: Transformers [9] excel in capturing global dependencies and enabling parallel computation. Innovations include sparse attention (Informer [10]), channel independence modeling (iTransformer [11]), frequency-domain enhancement (FEDformer [12]), and causality constraints (PathTST [13]). Challenges remain, including quadratic complexity with sequence length and limited explicit modeling of inherent time series attributes.
@@ 2.2 Time Series Foundation Models
Approach 1: Pre-training from Scratch. This approach trains models directly on large-scale time series data (e.g., sensor, financial, or weather data) to capture long-term dependencies, periodicity, and trends using self-supervised tasks like forecasting, imputation, or generation. These models are powerful for time-series tasks but require massive labeled data and high computational costs. Notable examples include TimeGPT [14], a Transformer-based model supporting zero-shot transfer; Tempo [15], a diffusion-based generative model for multivariate time series; TimesFM [16], a decoder-only architecture with hierarchical time aggregation for irregular data; and Moirai [17], a unified model with multi-task pretraining for long sequences.
Approach 2: Adapting LLMs for Time Series. This approach maps time series data into textual or other modalities using cross-modal alignment, leveraging LLMs’ semantic understanding for tasks like trend description or multimodal reasoning. Challenges include bridging the gap between continuous time series and discrete text, as well as handling causality conflicts with LLM position encodings. Key works include LLMTime [18], which discretizes time series into ”time tokens” for forecasting via text generation; TEST [19], which converts time series into natural language templates for zero-shot forecasting; TimeLLM [20], aligning time series with LLMs using embeddings; TEMPO [15], jointly encoding time series and text for domain-specific tasks; and Lag-Llama [21], fine-tuning Llama with lag operators for univariate forecasting.
@@ 3 Preliminaries
Transformer-based Models.   Transformer [9] consists of $L$ stacked blocks, each with multi-head attention (MHA) and feed-forward network (FFN) modules. For input $X\in\mathbb{R}^{n\times d}$, MHA computes:
@@
Residual connections and layer normalization follow both sub-modules. ReLU is often replaced by Gated Linear Units (GLU) [22]:
@@
Patching.   Time series forecasting benefits from treating subsequences (patches) as tokens rather than individual time points. This approach captures local semantic information and enhances predictive representations [13].
Channel Independence.   Multivariate time series can be modeled with channel independence, where each token represents data from a single channel [13]. This simplifies the model and allows learning channel-specific dynamics.
MOMENT Model.   MOMENT [23] is a Transformer-based time series foundation model pretrained on diverse data using masked prediction. It supports long-term/short-term prediction, anomaly detection, and classification. MOMENT-base consists of 12 Transformer encoders with $d=768$, $h=12$, and $d_{m}=3072$. Its FFN uses GLU:
@@ 4 Our method
4.1 Analyzing AdaLoRA’s solution
AdaLoRA parameterizes the incremental update of the pre-trained weight matrix $W$ using singular value decomposition:
$W=W^{(0)}+\Delta=W^{(0)}+P\Lambda Q,$ (6)
where $P\in\mathbb{R}^{d_{1}\times r}$ and $Q\in\mathbb{R}^{r\times d_{2}}$ represent the left and right singular vectors of $\Delta$, and $\Lambda\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing singular values $\{\lambda_{i}\}_{1\leq i\leq r}$.
The sensitivity of a parameter $w_{ij}$ is defined as:
$I(w_{ij})=|w_{ij}\nabla_{w_{ij}}\mathcal{L}|,$ (7)
where $\mathcal{L}$ is the loss function. This measures the impact of $w_{ij}$ on the model’s performance, with larger values indicating higher importance.
For the $k$-th incremental matrix, AdaLoRA computes the importance of the $i$-th singular value and its corresponding singular vectors as:
$S_{k,i}=s(\lambda_{k,i})+\frac{1}{d_{1}}\sum_{j=1}^{d_{1}}s(P_{k,ji})+\frac{1}{d_{2}}\sum_{j=1}^{d_{2}}s(Q_{k,ij}),$ (8)
where $s(\cdot)$ is a function of $I(\cdot)$. Singular values with lower importance are pruned by setting them to zero, effectively reducing the rank $\gamma$ of each incremental matrix.
Despite its strengths, AdaLoRA faces notable challenges in time-series scenarios. For instance, its SVD-based pruning prioritizes singular values reflecting global feature importance, potentially overlooking critical local signals such as abrupt anomalies or short-term seasonality. Furthermore, continuous rank adjustment introduces redundant parameters for low-dimensional, localized time-series embeddings, contradicting the principle of minimal fine-tuning. Finally, AdaLoRA’s static pruning is irreversible, risking underfitting if early pruning discards modules relevant to later training stages—a limitation addressed by Gated DSIC, which enables dynamic reactivation via differentiable gates.
4.2 Advantages of TRACE Over AdaLoRA
TRACE offers several key advantages over AdaLoRA, particularly in temporal contexts. First, Gated DSIC enforces conditional parameter consistency through a gating mechanism, ensuring unbiased importance estimation between masked and unmasked LoRA modules. This prevents irreversible loss of temporal features, such as recovering suppressed modules when seasonal patterns re-emerge. Additionally, the gate prioritizes LoRA modules aligned with local temporal dependencies, validated by ablation studies, whereas AdaLoRA’s global SVD may dilute transient signals. Importantly, Gated DSIC achieves comparable accuracy to full LoRA with 30–50% fewer active parameters, making it ideal for deploying lightweight models on edge devices.
Beyond dynamic gating, TRACE also introduces a parameter-efficient forecasting head reconstruction. Linear probing heads scale poorly with prediction horizons due to quadratic growth, but our techniques (e.g., 2D convolution, pooling) decouple head size from horizon length, reducing parameters by 4–8 times while maintaining accuracy. Furthermore, while AdaLoRA focuses solely on backbone adaptation, TRACE jointly optimizes sparse backbone updates (via Gated DSIC) and compact heads, mitigating overfitting on limited temporal data.
Empirical evaluations across 12 time-series benchmarks confirm TRACE’s superiority. In long-term forecasting, TRACE outperforms AdaLoRA by 2.1–3.8% in MSE using 30% fewer parameters. In few-shot anomaly detection, Gated DSIC achieves higher parameter efficiency, yielding 0.5–2.1% accuracy gains where localized adaptation is critical. Notably, Gated DISC matches AdaLoRA’s performance on the GLUE benchmark, demonstrating its generalizability beyond time-series.
@@ 4.2 Advantages of TRACE Over AdaLoRA
Figure 1: architecture of our method:(a) Represents the default model architecture, which consists of multiple stacked Transformer encoders. LoRA modules are added to all linear layers at each level, and the prediction head is a linear predictor. Fine-tuning is performed using linear probing and full LoRA adaptation.(b) Represents the TRACE method: LoRA integration follows the DSIC approach, and for long-term forecasting tasks, the prediction head is dimensionally reduced and reconstructed.
@@ 4.3 Reconstruction of Long-horizon Forecasting
MOMENT utilizes lightweight prediction heads, enabling fine-tuning with a limited number of trainable parameters. For long-term forecasting tasks, however, the number of parameters in the prediction head can become excessively large. The number of parameters is given by $H\times N\times d$, where $N$ is the number of patches ($N=64$), $d$ is the model dimension ($d=768$), and $H$ is the prediction horizon. For example, in MOMENT-base, with a prediction length of 192, the number of parameters is 9,437,184; for a prediction length of 720, it increases to 35,389,440. While this is relatively small compared to the total 125 million parameters of the entire MOMENT-base model, it is still significant when fine-tuning for downstream tasks. To address this, we focus on two strategies to reduce parameters:
(1) Reducing the representation dimension $d$ of each patch.
(2) Reducing the number of parameters associated with the prediction horizon $H$.
For the long-term prediction head $W_{(N\times d)\times H}$, we implement the following reconstruction methods:
@@
[…]
To reduce the dimensionality of patch representations, we experiment with several techniques: projection down, feature truncation, average pooling, and convolution. The patch dimension is reduced to $1/\beta$ of the original 768 dimensions, where $\beta=4,8,16,\dots$, referred to as the dimensionality reduction factor.
In subsequent experiments, we demonstrate that most reconstructed prediction heads perform equally well or better than the original MOMENT prediction head when fine-tuned on downstream time series data. Table 1 presents the reduction in the number of parameters and the corresponding reduction ratios for different values of $\beta$.
Table 1: Comparison of prediction head parameters across different horizons and reduction under varying $\beta$ values
Horizon $\beta$ Head Parameters Reduced Reduction Percentage
192 4 9437184 2506752 73.4%
8 9437184 1253376 86.7%
16 9437184 626688 93.4%
720 4 35389440 8994816 74.6%
8 35389440 4497408 87.3%
16 35389440 626688 98.2%
A is a concise description of the reconstruction operations applied to the forecast head. The parameter $d$ represents the embedding dimension for each patch (subsequence), with $d=768$ in MOMENT. To leverage the pre-trained model’s predictive capabilities and reduce overfitting, we reduce $d$ since patch length limits information content.
4.4 Our Gated DSIC method
4.4.1 Our solution setting
We define the LoRA Gate as follows:
$g_{ij},||g_{ij}||=1$ (11)
where $i$ represents the $i$-th layer, and $j$ represents the $j$-th LoRA module in the $i$-th layer. The LoRA Gate acts as a $switch$ controlling whether the incremental matrix is activated. For the $i$-th layer and $j$-th pre-trained linear module $h_{ij}$, we add a low-rank incremental matrix:
$h_{ij}=W_{ij}^{(0)}x+g_{ij}\cdot\Delta x_{ij}=W_{ij}^{(0)}x+g_{ij}\cdot BAx,$ (12)
where $W^{(0)},\Delta\in\mathbb{R}^{d_{1}\times d_{2}},A\in\mathbb{R}^{r\times d_{2}}$, and $B\in\mathbb{R}^{d_{1}\times r}$, with $r\ll\{d_{1},d_{2}\}$. Since the backpropagation and parameter updates for the entire LoRA module are managed by $g_{ij}$, the importance of $g_{ij}$ reflects the contribution of the corresponding LoRA module to the overall model loss. Similarly, the importance of each LoRA Gate can be computed as:
$I(g_{ij})=|g_{ij}\nabla_{g_{ij}}\mathcal{L}|=|g_{ij}||\nabla_{g_{ij}}\mathcal{L}|=|\nabla_{g_{ij}}\mathcal{L}|$ (13)
With this method, we can quickly and accurately assess the contribution of each LoRA module without having to calculate the importance of each incremental matrix. Taking MOMENT as an example, there are 12 Encoder layers in total, and each layer contains 7 linear layers for attention heads and FFN. Thus, there are $12\times 7=84$ LoRA Gates, which we denote as $\mathcal{G}=\{g_{ij}|i=1,2,\dots,12;j=1,2,\dots,7\}$. We can simply control whether each LoRA module is activated by setting $g_{ij}$ to 0 or 1.
@@ 4.4.1 Our solution setting
Our approach uses ”the sum of the gradient magnitudes of the unmasked gates as the importance score”:
@@
$I(g_{kl}|\;g_{ij}=1,g_{ij}\in\mathcal{G})\neq I(g_{kl}|\;g_{mn}=0,g_{mn}\in\tilde{\mathcal{G}}\;and\;g_{pq}=1,g_{pq}\notin\tilde{\mathcal{G}})$ (15)

### v3

@@ Abstract
We propose an efficient fine-tuning method for time series foundation models, termed TRACE: Time Series Parameter Efficient Fine-tuning. While pretrained time series foundation models are gaining popularity, they face the following challenges: (1) Time series data exhibit significant heterogeneity in frequency, channel count, and sequence lengths, necessitating tailored fine-tuning strategies, especially for long-term forecasting.(2) Existing parameter-efficient fine-tuning (PEFT) methods, such as LoRA, are not directly optimized for the unique characteristics of time series data.
To address these challenges, our TRACE framework introduces two key innovations: (1) Gated DSIC (Gated Dynamic Simulation Importance Calculation), a less biased LoRA module selection mechanism that ensures conditional parameter consistency before and after masking; and (2) Reconstructed forecasting heads for long-term forecasting tasks, which achieve comparable or superior performance to linear probing while drastically reducing parameter counts.
@@ 1 Introduction
Pre-trained time series foundation models (e.g., MOMENT [moment]) have demonstrated strong generalization across diverse forecasting and anomaly detection tasks. However, their effective adaptation to downstream applications remains hindered by two key challenges in fine-tuning.
First, standard full fine-tuning is computationally expensive and often infeasible on resource-constrained hardware. In contrast, linear probing—freezing the backbone and tuning only the forecasting head—is efficient but suboptimal: it fails to exploit the rich temporal representations encoded in intermediate layers. This limitation is especially severe in long-horizon forecasting, where the forecasting head itself can contain millions of parameters (e.g., >35>35M for a 720-step horizon in MOMENT), leading to overfitting and degraded performance.
Second, existing parameter-efficient fine-tuning (PEFT) methods like LoRA were designed for natural language tasks and do not account for the unique structure of time series data. Naive application of such methods suffers from biased importance estimation during module selection, resulting in suboptimal adaptation.
To address these issues, we propose TRACE (Time Series Parameter Efficient Fine-tuning), a novel framework with two core components:
(1) Reconstructed forecasting heads that factorize the original forecasting layer, reducing its parameter count by over 70% while improving or matching performance.
(2) Gated DSIC (Gated Dynamic Simulation Importance Calculation), a debiased LoRA module selection mechanism that simulates post-pruning contexts via Monte Carlo masking to accurately estimate module importance.
Our contributions are summarized as follows:
(1) We design lightweight, high-performance forecasting heads tailored for long-horizon time series tasks.
(2) We introduce Gated DSIC, a novel PEFT strategy that mitigates evaluation bias and enables dynamic, reversible LoRA pruning.
(3) We conduct extensive experiments across long/short-term forecasting, anomaly detection, and even NLP benchmarks, showing that TRACE consistently outperforms full fine-tuning, linear probing, and prior PEFT baselines.
(4) We demonstrate that our method generalizes beyond time series, achieving state-of-the-art results on GLUE and QA tasks with minimal trainable parameters.
@@ 2.1 Time Series Modeling
Statistical Methods: Early methods like Moving Average (MA) and Exponential Smoothing [hyndman2002state] smooth trends but fail to capture complex dynamics. ARIMA [arima] improves accuracy by integrating Autoregressive (AR) and MA components but struggles with nonlinear relationships. VAR [var] models multivariate relationships but suffers from dimensionality issues. These methods are limited to univariate modeling and cannot effectively use covariates.
Machine Learning Methods: Gradient boosting trees (e.g., XGBoost [chen2016xgboost], LightGBM [ke2017lightgbm]) leverage feature engineering strategies, such as lag features and sliding window statistics, to model nonlinear patterns. While outperforming statistical methods, their performance depends heavily on feature quality and fixed window lengths limit long-term dependency modeling.
Deep Learning Methods: RNNs and variants like LSTM [lstm] and GRU [gru] capture temporal dependencies but suffer from inefficiency and vanishing gradients for long-range dependencies. TCNs [tcn] expand receptive fields via dilated convolutions but are constrained by fixed kernels.
Transformer Architectures: Transformers [vaswani2017attention] excel in capturing global dependencies and enabling parallel computation. Innovations include sparse attention (Informer [zhou2021informer]), channel independence modeling (iTransformer [liu2023itransformer]), frequency-domain enhancement (FEDformer [fedformer]), and causality constraints (PathTST [patchtst]). Challenges remain, including quadratic complexity with sequence length and limited explicit modeling of inherent time series attributes.
@@ 2.2 Time Series Foundation Models
Approach 1: Pre-training from Scratch. This approach trains models directly on large-scale time series data (e.g., sensor, financial, or weather data) to capture long-term dependencies, periodicity, and trends using self-supervised tasks like forecasting, imputation, or generation. These models are powerful for time-series tasks but require massive labeled data and high computational costs. Notable examples include TimeGPT [garza2023timegpt], a Transformer-based model supporting zero-shot transfer; Tempo [cao2023tempo], a diffusion-based generative model for multivariate time series; TimesFM [timesFM], a decoder-only architecture with hierarchical time aggregation for irregular data; and Moirai [moirai], a unified model with multi-task pretraining for long sequences.
Approach 2: Adapting LLMs for Time Series. This approach maps time series data into textual or other modalities using cross-modal alignment, leveraging LLMs’ semantic understanding for tasks like trend description or multimodal reasoning. Challenges include bridging the gap between continuous time series and discrete text, as well as handling causality conflicts with LLM position encodings. Key works include LLMTime [llmTimes], which discretizes time series into ”time tokens” for forecasting via text generation; TEST [sun2023test], which converts time series into natural language templates for zero-shot forecasting; TimeLLM [jin2023time], aligning time series with LLMs using embeddings; TEMPO [cao2023tempo], jointly encoding time series and text for domain-specific tasks; and Lag-Llama [lagLlama], fine-tuning Llama with lag operators for univariate forecasting.
@@ 3 Preliminaries
Transformer-based Models.   Transformer [vaswani2017attention] consists of $L$ stacked blocks, each with multi-head attention (MHA) and feed-forward network (FFN) modules. For input $X\in\mathbb{R}^{n\times d}$, MHA computes:
@@
Residual connections and layer normalization follow both sub-modules. ReLU is often replaced by Gated Linear Units (GLU) [shazeer2020glu]:
@@
Patching.   Time series forecasting benefits from treating subsequences (patches) as tokens rather than individual time points. This approach captures local semantic information and enhances predictive representations [patchtst].
Channel Independence.   Multivariate time series can be modeled with channel independence, where each token represents data from a single channel [patchtst]. This simplifies the model and allows learning channel-specific dynamics.
MOMENT Model.   MOMENT [moment] is a Transformer-based time series foundation model pretrained on diverse data using masked prediction. It supports long-term/short-term prediction, anomaly detection, and classification. MOMENT-base consists of 12 Transformer encoders with $d=768$, $h=12$, and $d_{m}=3072$. Its FFN uses GLU:
@@ 4 Our method
4.1 Gated DSIC: Mitigating Bias in LoRA Module
Parameter-efficient fine-tuning (PEFT) methods like LoRA offer a promising alternative to full fine-tuning. However, naively applying them to time series foundation models is suboptimal. For instance, AdaLoRA parameterizes the weight update via singular value decomposition (SVD) and prunes low-importance singular values to adapt the rank dynamically. While effective in some domains, AdaLoRA faces notable challenges in time-series scenarios. Its SVD-based pruning prioritizes singular values reflecting global feature importance, potentially overlooking critical local signals such as abrupt anomalies or short-term seasonality. Furthermore, its static, irreversible pruning risks underfitting if early-stage pruning discards modules that become relevant later in training.
To address these limitations, we propose Gated Dynamic Simulation Importance Calculation (Gated DSIC), a less biased LoRA module selection mechanism. Gated DSIC introduces a learnable gate for each LoRA module, enabling dynamic and reversible masking. Crucially, it employs a Monte Carlo simulation during the importance evaluation phase to ensure that the importance score of a module is calculated under conditions that mimic the post-masking state, thereby mitigating the bias inherent in sequential, deterministic pruning strategies like AdaLoRA. This ensures conditional parameter consistency before and after masking, allowing the model to recover suppressed modules if needed (e.g., when seasonal patterns re-emerge). Concretely, each LoRA module is associated with a gate whose activation is determined by a less biased importance score estimated via repeated random masking on the validation set. The detailed algorithm is presented in Section 4.3.
@@ 4.1 Gated DSIC: Mitigating Bias in LoRA Module
Figure 1: architecture of our method:(a) Represents the default model architecture, which consists of multiple stacked Transformer encoders. LoRA modules are added to all linear layers at each level, and the forecasting head is a linear predictor. Fine-tuning is performed using linear probing and full LoRA adaptation.(b) Represents the TRACE method: LoRA integration follows the Gated DSIC approach, and for long-term forecasting tasks, the forecasting head is dimensionally reduced and reconstructed.
@@ 4.2 Reconstruction of Long-horizon Forecasting
The standard linear forecasting head in models like MOMENT, parameterized by a matrix $\mathbf{W}\in\mathbb{R}^{(N\times d)\times H}$, becomes prohibitively large for long-horizon forecasting (e.g., $H=720$), often exceeding 35M parameters. While small relative to the full backbone, this introduces significant overfitting risk and computational burden during fine-tuning.
To address this, we propose reconstructing the forecasting head via low-rank factorization, reducing its parameter count while preserving or enhancing predictive performance. As illustrated in Figure 2, our approach decomposes the original weight matrix $\mathbf{W}$ into two smaller matrices:
$\mathbf{W}_{\text{recon}}=\mathbf{W}_{1}\cdot\mathbf{W}_{2},\quad\text{where}\quad\mathbf{W}_{1}\in\mathbb{R}^{(N\times d)\times d^{\prime}},\;\mathbf{W}_{2}\in\mathbb{R}^{d^{\prime}\times H},$ (6)
and $d^{\prime}=d/\beta$ is the reduced embedding dimension, controlled by the reduction factor $\beta$.
This reconstruction reduces the parameter count from $\mathcal{O}(NdH)$ to $\mathcal{O}(Nd^{2}/\beta+dH/\beta)$, achieving reductions of up to 98.2% (see Table 1) with minimal performance loss. We experiment with four variants—Projection Down (ProjDown), Feature Truncation (LessFeature), Average Pooling (AvgPool), and 2D Convolution (Conv2D)—each offering different trade-offs between efficiency and representational capacity.
Our experiments (Section 5.5.2) show that ProjDown and AvgPool consistently outperform the baseline linear probing, validating the effectiveness of our reconstruction strategy.
Figure 2: Illustration of forecasting head reconstruction. The original large matrix $\mathbf{W}\in\mathbb{R}^{(N\times d)\times H}$ is factorized into two smaller matrices, reducing the embedding dimension from $d$ to $d^{\prime}=d/\beta$. This significantly cuts the number of trainable parameters without sacrificing predictive power.
Table 1: Comparison of forecasting head parameters across different horizons and reduction under varying $\beta$ values
Horizon $\beta$ Head Parameters Reduced Reduction Percentage
192 4 9,437,184 2,506,752 73.4%
8 9,437,184 1,253,376 86.7%
16 9,437,184 626,688 93.4%
720 4 35,389,440 8,994,816 74.6%
8 35,389,440 4,497,408 87.3%
16 35,389,440 626,688 98.2%
A is a concise description of the reconstruction operations applied to the forecasting head. The parameter $d$ represents the embedding dimension for each patch (subsequence), with $d=768$ in MOMENT. To leverage the pre-trained model’s predictive capabilities and reduce overfitting, we reduce $d$ since patch length limits information content.
4.3 Our Gated DSIC method
4.3.1 Learnable Gating for Dynamic LoRA Control
To enable dynamic and reversible selection of LoRA modules during parameter-efficient fine-tuning, we introduce a learnable scalar gate $g_{ij}\in\mathbb{R}$ for the $j$-th LoRA module in the $i$-th Transformer layer. This gate multiplicatively modulates the low-rank update, yielding the adapted layer output:
$h_{ij}=W^{(0)}_{ij}x+g_{ij}\cdot\Delta x_{ij}=W^{(0)}_{ij}x+g_{ij}\cdot BAx,$
where $W^{(0)}_{ij}$ denotes the pre-trained weight matrix, and $A\in\mathbb{R}^{r\times d_{2}}$, $B\in\mathbb{R}^{d_{1}\times r}$ are the low-rank adaptation matrices with rank $r\ll\min(d_{1},d_{2})$. The gate $g_{ij}$ is initialized to 1 and optimized jointly with $A$ and $B$ via backpropagation.
Crucially, the importance of each LoRA module is estimated through the magnitude of the gradient of the validation loss with respect to its gate:
@@ 4.3.1 Learnable Gating for Dynamic LoRA Control
This gradient reflects how sensitive the model’s performance is to perturbations in $g_{ij}$, providing a natural proxy for module contribution.
Taking MOMENT-base as an example, which consists of 12 Transformer encoder layers with 7 linear sub-layers each (Query, Key, Value, Output, and three FFN projections), this yields a total of 84 gates, denoted as $\mathcal{G}=\{g_{ij}\mid i=1,\dots,12;\,j=1,\dots,7\}$.
@@
Following common practice, we define the importance score of unmasked gates as the expected sum of gradient magnitudes:
@@
[…]